OpenStudy (anonymous):

who can help me to Solve this problem. Brandon can paint a fence in 12 hours and Elaine can paint the same fence in 11 hours. How long will they take to paint the fence if they work together? Please choose the rihgt answer. a. 5 \frac{13}{24} hr b. 5 \frac{1}{4} hr c. 5 \frac{17}{23} hr d. 11 \frac{1}{2} hr

OpenStudy (anonymous):

\[\frac{11\times 12}{11+12}\]

OpenStudy (anonymous):

a. 5 \frac{13}{24} hr b. 5 \frac{1}{4} hr c. 5 \frac{17}{23} hr d. 11 \frac{1}{2} hr Which is the right answer?

OpenStudy (anonymous):

Hi who knows this question?